Please help me answer the following question: If 2x-3(x-4)=-5, then x=?
Answer by checkley79(3341)

2x-3(x-4)=-5
2x-3x+12=-5
-x=5-12
-x=-7 x=7 ans.
Proof:
2*7-3(7-4)=5
14-3*3=5
14-9=5
5=5